OpenStudy (anonymous):

From 4a + 7b - 12c subtract 8a - 9b + 12d. I need help .

OpenStudy (anonymous):

this is a tricky problem to start, but once you undertand it, it will be a lot easier! we have to interprete the words into math symbols From 4a + 7b - 12c subtract 8a - 9b + 12d. From 4a + 7b - 12c - (8a - 9b + 12d) do you see how I did that?

OpenStudy (anonymous):

so how would i get my answer

OpenStudy (anonymous):

you do it for each variable

OpenStudy (anonymous):

we have this, and then we 'distribute' the '-' sign 4a + 7b - 12c - (8a - 9b + 12d) 4a + 7b - 12c - 8a + 9b - 12d

OpenStudy (anonymous):

i had -4a + 16b - 24c - 24d but it was partial credit

OpenStudy (anonymous):

4a-8a=-4a 7b-(-9b)= 7b+9b=16b (-12c)-12c=-24c -4a+16b-24c

OpenStudy (anonymous):

what about for d?

OpenStudy (anonymous):

do you see how you ended up with a "- 24c - 24d" it should just be "- 12c - 12d"

OpenStudy (anonymous):

so the answer is -4a+16b -12c - 12d?

OpenStudy (anonymous):

yep, becuase we can only combine 'like terms'

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

^_^